OpenStudy (anonymous):

I need help right now. My problem is 8/9 - 14/27. Suppossed to lower the fraction down. Can you help?

OpenStudy (katrinakaif):

"Lower the fraction?"

OpenStudy (anonymous):

24/27 = 8/9 24/27 - 14/27 = 10/27

OpenStudy (anonymous):

10/27 is lowest form

OpenStudy (katrinakaif):

Find the common denominator between both of them. Here, 9 x 3 = 27 giving us similar ones between both fractions. So you must multiply x 3 on the top as well, resulting 8 x 3 = 24 Your first fraction ----> 24/27 Now substract: 24/27 - 14/27 = 10/27 =) It cannot be further simplified, thus resulting to be your answer.